# MDCure: A Scalable Pipeline for Multi-Document Instruction-Following

### v1

@@ 5.2 Influence of Data Scale
None 4.4 45.1 1.74 2.16 1.7 13.1
12K RM-Filtered 43.8 48.4 4.16 2.43 1.9 22.0
36K RM-Filtered 45.7 46.1 4.17 2.55 1.9 22.2
72K RM-Filtered 47.3 48.3 4.22 2.58 2.1 22.6
72K Unfiltered 31.7 47.5 4.04 2.35 1.9 21.0
@@
None 30.5 39.6 4.30 3.57 7.4 23.9
12K RM-Filtered 44.1 45.4 4.26 3.72 8.9 25.9
36K RM-Filtered 44.4 45.8 4.26 3.81 8.6 28.1
72K RM-Filtered 44.7 46.0 4.28 3.93 10.3 29.8
72K Unfiltered 40.5 43.3 4.26 3.70 8.5 26.3
@@
In order to understand the importance of data quantity, we compare the impact of fine-tuning with MDCure instruction datasets of various sizes. We experiment with FlanT5-Base and Qwen2-7B-Instruct as the base models to demonstrate the effect of data size at different model scales. Results are compared versus the non-MDCure’d baselines and shown in Table 3. We observe consistent, albeit modest, improvements to downstream MD performance as we scale up the instruction data size. Notably, use of 12K MDCure-filtered samples consistently surpasses performance with use of 72K unfiltered samples for both base models, consistent with other work finding that use of only a few thousand high-quality instruction samples is sufficient for alignment (Zhou et al., 2023). This evidences MDCureRM’s ability to promote data efficiency and reduce costs involved in generating and using an effective MD instruction-tuning dataset.
5.3 Generalization to Few-Shot & Long-Context Training
We investigate the scalability of MDCure-generated instructions to the few-shot prompting and long-context settings. The few-shot setting is of interest as it provides an inexpensive alternative for adapting models to new tasks with limited compute (Dong et al., 2023). Likewise, long-context processing is important for real-world tasks over “extreme” numbers of documents such as extreme MDS (Lu et al., 2020). We extend MDCure instruction data to these settings either by appending distractor articles to existing MDCure instruction inputs or by packing few-shot exemplars into the input context up to a context length of 32k tokens. [^7] Context length selected based on GPU memory limits. Distractor articles are identified by embedding all NewSHead documents using the Sentence Transformers model all-distilroberta-v1, [^8] We choose this embedding model from https://www.sbert.net/docs/sentence_transformer/pretrained_models.html due to its strong performance, long context length, and compatibility with cosine similarity. and taking the most cosine-similar documents to the documents in the source cluster. We choose similar documents instead of performing random selection to increase the difficulty of the instruction task and encourage identification of relevant documents in a long context, and since recent work finds using sequences of related documents during training can assist models to read and reason across document boundaries (Shi et al., 2023). For few-shot training, exemplars are chosen randomly, with each exemplar appearing once across the duration of training.
@@ 5.3 Generalization to Few-Shot & Long-Context Training
LLAMA3.1-8B-Instruct 35.5 27.1 4.28 2.94 10.2 18.7
+MDCure 41.0 40.2 4.30 3.94 12.0 31.7
+MDCure +LongContext 38.9 40.5 4.31 3.97 12.1 32.5
@@
LLAMA3.1-8B-Instruct 57.5 44.2 3.97 2.93 15.1
+MDCure 63.9 47.0 4.04 3.32 17.3
+MDCure +Few-Shot 65.0 49.2 4.08 3.32 17.9
@@
We conduct experiments with LLAMA3.1-8B-Instruct as the base model, with details of the training setup described in App. B.1. In the long-context setting, we use the same zero-shot prompting evaluations as before, comparing model performance when trained using distractor-extended instruction data versus when trained with non-extended instruction data and versus the non-MDCure’d baseline. In the few-shot setting, we perform similar comparisons but utilize 5-shot prompting during evaluation, leaving out ZeroScrolls since it is built for the zero-shot setting. Results are displayed in Table 4. We find that combining MDCure instruction data with longer contexts improves performance for most benchmarks. Similarly, incorporating multiple MDCure datapoints together to construct few-shot training inputs yields consistent improvements in few-shot evaluation. This confirms that our proposed instruction generation pipeline is extensible and beneficial for long-context tasks.
5.4 Effect of Generation Approach
@@
The efficacy of MDCure is dependent on the quality of initial candidate instruction generations. To this end, we investigate the impact of several different prompting approaches on downstream instruction-tuned model performance. We consider only the zero-shot prompting setting given the high cost of prompting GPT models with few-shot examples that each contain multiple long documents.
@@
In our preliminary investigation, we find that prompt templates emphasizing the MD nature of generated instructions tend to yield higher-scoring results (e.g., suggesting penalties if generated instructions concern only a single document). Thereafter, we explore the utility of providing general vs. highly detailed specifications for the style of the generated instruction-answer pairs. We ultimately converge on two classes of prompt templates which we refer to as either General or Style-Specific. Details are given in in App. A.2.
@@
Baseline (No IT) 4.4 45.1 1.74 2.16 1.7 13.1
General 42.3 40.5 4.09 2.34 1.1 20.6
Style-Specific 37.9 48.1 4.13 2.20 1.3 20.0
Both 43.8 48.4 4.16 2.43 1.9 22.0
@@
Baseline (No IT) 30.5 39.6 4.30 3.57 7.4 23.9
General 40.8 44.9 4.22 3.66 7.8 24.2
Style-Specific 40.2 45.8 4.24 3.63 7.8 23.2
Both 44.1 45.4 4.26 3.72 8.9 25.9
@@
We examine the impact of General vs. Style-Specific prompt templates by fine-tuning FlanT5-Base and Qwen2-7B-Instruct using MDCureRM-filtered datasets created using templates from one or both categories. We limit the number of samples to 12K since §5.2 finds this quantity to be sufficient for strong improvements to MD task performance. When both templates are used, we employ a 1:3 ratio of General:Style-Specific instructions, found to be most effective in initial experiments.
5.5 Ablations on Scoring Criteria
To justify our choice of scoring criteria in MDCureRM, we conduct an ablation study to show the advantage of combining all six criteria, rather than using a subset of them. We also evidence the benefit of applying greater weight to MD-specific criteria when computing the overall score for each candidate instruction, as opposed to computing an evenly weighted sum (details in App. A.6). Ablation experiments were run using FlanT5-Base and Qwen2-7B-Instruct. Table 6 shows the results. Across all benchmarks, filtering instruction data using a weighted average of six scoring criteria with emphasis on multi-document factors yields the best results. Further, removing any one criterion or utilizing even weighting across criteria leads to worsened IT’ed performance.
@@ 5.5 Ablations on Scoring Criteria
all criteria, MD emphasis 43.8 48.4 4.16 2.43 1.9 22.0
all criteria, evenly weighted 43.4 47.1 4.16 2.35 1.8 21.2
without Relevance 37.6 46.1 4.12 2.26 1.6 20.4
without Coh. & Fac. 40.1 46.0 4.14 2.35 1.5 20.4
without Creativity 38.4 44.2 4.15 2.33 1.3 20.8
without Context Int. 14.1 45.4 3.84 2.06 1.5 17.0
without Inter-Doc. Rel. 40.9 45.8 4.10 2.41 1.5 21.5
without Complexity 40.2 45.5 4.11 2.44 1.4 20.0
@@
all criteria, MD emphasis 44.1 45.4 4.26 3.72 8.9 25.9
all criteria, evenly weighted 41.3 43.1 3.97 3.10 8.5 25.4
without Relevance 37.5 40.7 4.25 3.69 8.1 25.0
without Coh. & Fac. 41.9 43.8 4.27 3.63 8.4 23.2
without Creativity 43.6 44.3 4.23 3.72 8.9 25.4
without Context Int. 37.1 42.4 4.24 3.62 8.5 25.0
without Inter-Doc. Rel. 39.0 42.4 4.24 3.65 8.5 25.7
without Complexity 37.1 41.8 4.25 3.65 8.5 25.7
@@
6 Conclusion
In this work, we presented MDCure, a procedure for automatically generating high-quality multi-document instruction tuning data to improve any base LLM’s ability to leverage long range dependencies for MD tasks. MDCure addresses the challenge of data scarcity in MD task domains and is general and scalable across model families and sizes. We further introduced MDCureRM, a novel RM-as-a-Judge LLM that can be used to filter MD instruction data more cheaply and effectively than GPT-3.5-Turbo. Our extensive experiments on various LLMs in multiple task domains demonstrate that MDCure adds value beyond typical instruction tuning and can substantially boost MD performance while overcoming limitations of previous pre-training-based approaches. Future work could explore extending MDCure to other topic domains, or use of MDCureRM to train a custom MD instruction generator model.
7 Limitations
While we have obtained promising experimental results, we recognize several limitations of this work. In our experiments we limit generation of MD instruction data to the news domain to preserve comparability to prior state-of-the-art MD pre-training approaches. Applying MDCure to document sets in other domains and text genres (e.g., science, mathematics) may yield further boosts to instruction diversity and generalization performance of resulting IT’ed models. While MDCure is able to accommodate use of arbitrarily large document collections for instruction generation, our experiments are limited to generation over 3-5 articles due to limited cluster size in the NewSHead source dataset. Scaling up the number or length of context documents may require use of more powerful and costly generator LLMs such as GPT4 to ensure sufficient quality of the raw generations, which place an upper bound on the overall composition of any possible resulting instruction set. This may limit the scalability of the approach; however, combined with the costs saved through use of our custom RM-as-a-Judge model, we believe MDCure provides a promising avenue to more effectively balance costs associated with creating high-quality instruction data versus existing synthetic data pipelines.
Experimentally, our training is limited to 4096 context length due to compute restrictions. Given the results in §5.2 and §5.3, we expect that our MDCure’d models could achieve greater performance on MD benchmarks if more instruction samples and greater training context length were used. The design and application of our approach is limited to texts in English, suggesting limited utility of MDCure for non-English tasks. Further, despite our exploration of a wide range of prompts for raw instruction generation, it is plausible that slight changes in phrasing could still affect the quality of resulting generations, and different generator models may need different prompts to achieve similar effect. Potential improvements may arise from generating instructions and answers separately as opposed to together (Zhao et al., 2024a), and from incorporation of short-context data during instruction-tuning (Bai et al., 2024; Gao et al., 2024b). Nonetheless, instruction complexity and answer correctness are constrained by the ability for the generator model to perform such MD tasks.
8 Ethics Statement
As MDCure creates instruction data based on real-world text corpora, one potential consequence is amplification of biases from the source texts in resulting IT’ed LLMs. Source texts are generally not filtered prior to use in MDCure, so it is possible to enable creation of instruction tuning samples concerning unethical or sensitive content. As with any use of LLMs, there is a need to ensure safety of IT’ed models’ responses through incorporation of red-teaming or other safety measures. Use of our factuality criterion in assessing instruction data reduces the chances for MDCure to lead to unfactual examples due to hallucination during instruction generation. However, faithfulness to false anecdotal information in source texts could take precedence, leading to propagation of potentially biased content. More beneficially, we anticipate our work will improve LLMs’ efficacy in processing long documents or large document collections, which can accelerate value generation in domains such as business analytics and scientific literature mining.
Acknowledgments
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-2139841, We are grateful for the compute support provided by Google. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or Google.
Appendix A Dataset Construction
A.1 Source Texts

### v3

@@ 5.2 Influence of Data Scale
None 4.4 45.1 38.7 48.0 1.7 13.1
12K RM-Filtered 43.8 48.4 92.4 54.0 1.9 22.0
36K RM-Filtered 45.7 46.1 92.7 56.7 1.9 22.2
72K RM-Filtered 47.3 48.3 93.8 57.3 2.1 22.6
72K Unfiltered 31.7 47.5 89.8 52.2 1.9 21.0
@@
None 30.5 39.6 95.6 79.3 7.4 23.9
12K RM-Filtered 44.1 45.4 94.7 82.7 8.9 25.9
36K RM-Filtered 44.4 45.8 94.7 84.7 8.6 28.1
72K RM-Filtered 44.7 46.0 95.1 87.3 10.3 29.8
72K Unfiltered 40.5 43.3 94.7 82.2 8.5 26.3
@@
To understand the importance of data quantity, we compare the impact of fine-tuning with MDCure instruction datasets of various sizes. We utilize FlanT5-Base and Qwen2-7B-Instruct as base models to demonstrate the effect of data size at different model scales. Results are compared versus non-MDCure’d baselines and shown in Table 4. We observe consistent, albeit modest, improvements to downstream MD performance as we scale up the instruction data size. Notably, use of 12K MDCure-filtered samples consistently surpasses performance with use of 72K unfiltered samples for both base models, consistent with other work finding use of only a few thousand high-quality samples is sufficient for alignment (Zhou et al., 2023). This evidences MDCureRM’s ability to promote data efficiency and reduce costs involved in generating and using an effective MD IT dataset.
@@
LLAMA3.1-8B-Instruct 35.5 27.1 95.1 65.3 10.2 18.7
+MDCure 41.0 40.2 95.6 87.6 12.0 31.7
+MDCure +LongContext 38.9 40.5 95.8 88.2 12.1 32.5
@@
LLAMA3.1-8B-Instruct 57.5 44.2 88.2 65.1 15.1
+MDCure 63.9 47.0 89.8 73.8 17.3
+MDCure +Few-Shot 65.0 49.2 90.7 73.8 17.9
@@
5.3 Generalization to Few-Shot & Extended-Context Training
We investigate the scalability of MDCure instructions to the few-shot prompting and extended-context settings. The few-shot setting is of interest as it provides an inexpensive alternative for adapting models to new tasks with limited compute (Dong et al., 2023). Likewise, real-world MD tasks often occur over “extreme” numbers of documents such as in extreme MDS (Lu et al., 2020). We extend our MDCure data to these settings by packing few-shot exemplars or distractor articles into instruction inputs up to 32K tokens in length. [^8] Details of extended-context data construction is in App. B.11. Context length is selected based on GPU memory limits. Experiments utilize LLAMA3.1-8B-Instruct as the base model, with training setup details described in App. C.1. Results are displayed in Table 5, where we evaluate the few-shot setting via 5-shot prompting, omitting ZeroScrolls which targets 0-shot evaluation. We find that use of MDCure with extended contexts further improves performance for most benchmarks. Combined with prior results showing MDCure generalizes beyond training context length, this suggests MDCure is effective and applicable in both resource-rich and resource-constrained settings, enhancing its practicality.
5.4 Efficacy with Open-Source Generation
We demonstrate the distinct cost advantage of MDCure over existing synthetic data pipelines which rely on GPT-based generations and target only non-MD contexts by studying the impact of using an open-source generator model in place of GPT-3.5-Turbo. Following the same procedure as in previous settings, we alternately use LLAMA3.1-70B-Instruct generation with MDCureRM filtering to create an MDCure instruction dataset of size 12K examples, since §5.2 finds this quantity to be sufficient for strong improvements to MD task performance. Fine-tuning results on Qwen2-1.5B-Instruct, Qwen2-7B-Instruct, and LLAMA3.1-8B-Instruct are reported in Table 6. Across model sizes and families, use of open-source generations in fact leads to better performance versus use of GPT generations. These emphasize the flexibility of MDCure and validate the unique cost-saving nature of our synthetic data framework.
Generator LLM HQA WH MXS QC SEAM ZSS Avg
Qwen2-1.5B-Instruct
GPT-3.5-Turbo 37.0 33.2 94.4 83.3 10.0 24.4 42.3
LLAMA3.1-70B 36.7 33.9 94.2 82.4 10.3 24.5 42.8
Qwen2-7B-Instruct
GPT-3.5-Turbo 44.1 45.4 94.7 82.7 8.9 25.9 43.0
LLAMA3.1-70B 46.9 45.6 94.9 82.7 8.2 28.7 44.0
LLAMA3.1-8B-Instruct
GPT-3.5-Turbo 40.9 41.6 95.8 87.6 11.9 30.7 50.3
LLAMA3.1-70B 40.0 37.5 97.3 88.0 12.2 34.3 52.9
Table 6: MDCure results with different generator LLMs.
5.5 Integration with Policy Optimization
Observing MDCureRM’s strength as a judge of MD instruction quality, we explore the value of using MDCureRM with reinforcement learning methods such as PPO to iteratively improve synthetic MD data generations. We utilize LLAMA3.1-8B-Instruct as the initial policy model [^9] We select this model as it is relatively small while remaining effective at instruction-following. and use MDCureRM to provide reward signals whilst training with a PPO objective. Details of our PPO training setup can be found in App. C.3. As in §5.4, we use the final policy model to create a size-12K MDCure instruction dataset and report fine-tuning results on Qwen2-1.5B-Instruct, Qwen2-7B-Instruct, and LLAMA3.1-8B-Instruct in Table 7. Across benchmarks, use of generations from PPO-trained LLAMA3.1-8B-Instruct enables all three models to achieve close or superior performance versus use of GPT generations. Notably, MDCure enables models far smaller than SOTA LLMs to become effective MD instruction generators, further enhancing the cost-effectiveness of our approach.
Generator LLM HQA WH MXS QC SEAM ZSS Avg
Qwen2-1.5B-Instruct
GPT-3.5-Turbo 37.0 33.2 94.4 83.3 10.0 24.4 42.3
Llama3.1-8B+PPO 38.5 34.0 94.3 81.9 11.3 24.8 43.5
Qwen2-7B-Instruct
GPT-3.5-Turbo 44.1 45.4 94.7 82.7 8.9 25.9 43.0
Llama3.1-8B+PPO 46.7 47.8 95.1 82.5 9.1 26.9 45.3
LLAMA3.1-8B-Instruct
GPT-3.5-Turbo 40.9 41.6 95.8 87.6 11.9 30.7 50.3
Llama3.1-8B+PPO 41.5 40.6 94.9 93.5 11.8 32.1 53.1
Table 7: MDCure performance when MDCureRM is used to train a custom MD instruction generator model.
5.6 Impact on Advanced Model Capabilities
We assess the impact of MDCure on general LLM capabilities including language understanding, advanced reasoning, mathematics, and coding by evaluating our models at the 7B and 8B parameter scales on GSM8K (Cobbe et al., 2021b), MBPP (Austin et al., 2021), and MMLU (Hendrycks et al., 2021a, b) via the OLMES framework (Gu et al., 2024). Results are shown in Table 8. We observe that MDCure’d models achieve comparable average performance across the three benchmarks versus non-MDCure’d base models, demonstrating that MDCure data enhances MD modeling abilities of LLMs while preserving other advanced capabilities as well.
6 Conclusion
In this work, we presented MDCure, a framework for scalably and inexpensively generating high-quality MD post-training data to improve any base LLM’s ability to leverage long-range dependencies for MD tasks. MDCure addresses the challenge of data scarcity in MD task domains without depending on proprietary LLMs and is robust across model families and sizes. We further introduced MDCureRM, a novel RM-as-a-Judge LLM to filter MD instruction data more cheaply and effectively than GPT-3.5-Turbo and enable moderately-sized LLMs to excel over SOTA proprietary counterparts as generators of high-quality MD instruction data. Our extensive experiments on various LLMs in many task and content domains show MDCure adds value beyond typical IT and substantially boosts MD performance while overcoming limitations of pre-training-based approaches. Future work could explore extending MDCure to generate instructions over additional topic domains.
Model GSM8K MBPP MMLU Avg
Qwen2-7B-Instruct 0.781 0.634 0.706 0.707
+MDCure 0.801 0.668 0.694 0.721
LLAMA3.1-8B-Instruct 0.800 0.669 0.717 0.729
+MDCure 0.806 0.720 0.713 0.746
Table 8: General performance of MDCure’d models.
7 Limitations
As abstractive tasks are the focus of many downstream MD applications, MDCure focuses on instruction generation for abstractive settings; future work can consider incorporation of extractive instruction templates to expand MDCure’s utility. Additionally, the design and application of our approach is limited to texts in English; future work can also consider exploring the utility of MDCure for non-English tasks. As observed with existing state-of-the-art generative language models, post-trained models may still suffer from factuality errors during generation. Although the focus of this work is not on factuality, we adopt measures to improve factual consistency by eliminating non-factual or unfaithful instruction generations in the Filtering stage; faithfulness aspects of our pipeline could be the subject of other future investigation. Lastly, potential improvements may arise from generating instructions and answers separately as opposed to together (Zhao et al., 2024a), from incorporation of short-context data during instruction-tuning (Bai et al., 2024; Gao et al., 2024b), and with inclusion of document sets in other text genres. Such design choices are left as possibilities for future research given the cost-effectiveness, efficacy, and domain-general nature of our approach.
8 Ethics Statement
As with any use of LLMs, there is a need to ensure the safety of our post-trained models’ responses through red-teaming or other safety measures. Use of our factuality criterion in assessing instruction data reduces the chances for MDCure to lead to unfactual examples due to hallucination during instruction generation. However, it is generally difficult to guarantee factual consistency during generative language modeling without further fine-tuning. Beyond this, we anticipate our work will improve LLMs’ efficacy in processing long documents and large document collections, which can accelerate value generation in domains such as business analytics and scientific literature mining.
Acknowledgments
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-2139841. We are grateful for the compute support provided by Google. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or Google.
Appendix A Additional Ablation Results
A.1 Effect of Generation Approach
@@
The efficacy of MDCure is dependent on the quality of initial candidate instruction generations. To this end, we conduct an ablation study to investigate the impact of several different prompting approaches on downstream instruction-tuned model performance. We consider only the zero-shot prompting setting given the high cost of prompting GPT models with few-shot examples that each contain multiple long documents.
@@
In our preliminary investigation, we find that prompt templates emphasizing the MD nature of generated instructions tend to yield higher-scoring results (e.g., suggesting penalties if generated instructions concern only a single document). Thereafter, we explore the utility of providing general vs. highly detailed specifications for the style of the generated instruction-answer pairs. We ultimately converge on two classes of prompt templates which we refer to as either General or Style-Specific. Details are given in in App. B.2.
We examine the impact of General vs. Style-Specific prompt templates by fine-tuning FlanT5-Base and Qwen2-7B-Instruct using MDCureRM-filtered datasets created using templates from one or both categories. As in §5.4, we limit the number of samples to 12K given the findings in §5.2. When both templates are used, we employ a 1:3 ratio of General:Style-Specific instructions, found to be most effective in initial experiments.
@@
Baseline (No IT) 4.4 45.1 38.7 48.0 1.7 13.1
General 42.3 40.5 90.9 52.0 1.1 20.6
Style-Specific 37.9 48.1 91.8 48.9 1.3 20.0
Both 43.8 48.4 92.4 54.0 1.9 22.0
@@
Baseline (No IT) 30.5 39.6 95.6 79.3 7.4 23.9
General 40.8 44.9 93.8 81.3 7.8 24.2
Style-Specific 40.2 45.8 94.2 80.7 7.8 23.7
Both 44.1 45.4 94.7 82.7 8.9 25.9
@@
A.2 Ablations on Scoring Criteria
To justify our choice of scoring criteria in MDCureRM, we conduct an ablation study to show the advantage of combining all six criteria, rather than using a subset of them. We also evidence the benefit of applying greater weight to MD-specific criteria when computing the overall score for each candidate instruction, as opposed to computing an evenly weighted sum (details in App. B.9). Ablation experiments were run using FlanT5-Base and Qwen2-7B-Instruct. Table 10 shows the results. Across all benchmarks, filtering instruction data using a weighted average of six scoring criteria with emphasis on multi-document factors yields the best results. Further, removing any one criterion or utilizing even weighting across criteria leads to worsened IT’ed performance.
@@ A.2 Ablations on Scoring Criteria
all criteria, MD emphasis 43.8 48.4 92.4 54.0 1.9 22.0
all criteria, evenly weighted 43.4 47.1 92.4 52.2 1.8 21.2
without Relevance 37.6 46.1 91.6 50.2 1.6 20.4
without Coh. & Fac. 40.1 46.0 92.0 52.2 1.5 20.4
without Creativity 38.4 44.2 92.2 51.8 1.3 20.8
without Context Int. 14.1 45.4 85.3 45.8 1.5 17.0
without Inter-Doc. Rel. 40.9 45.8 91.1 53.6 1.5 21.5
without Complexity 40.2 45.5 91.3 54.2 1.4 20.0
@@
all criteria, MD emphasis 44.1 45.4 94.7 82.7 8.9 25.9
all criteria, evenly weighted 41.3 43.1 88.2 68.9 8.5 25.4
without Relevance 37.5 40.7 94.4 82.0 8.1 25.0
without Coh. & Fac. 41.9 43.8 94.9 80.7 8.4 23.2
without Creativity 43.6 44.3 94.0 82.7 8.9 25.4
without Context Int. 37.1 42.4 94.2 80.4 8.5 25.0
without Inter-Doc. Rel. 39.0 42.4 94.2 81.1 8.5 25.7
without Complexity 37.1 41.8 94.4 81.1 8.5 25.7
@@
Appendix B Dataset Construction
B.1 Source Texts